x+y=? x-7=5 x*y=24

Found 2 solutions by JimboP1977, ankor@dixie-net.com:
Answer by JimboP1977(311)

x-7=5
x=5+7
x=12
12*y=24
y=24/12
y=2
12+2=14

Answer by ankor@dixie-net.com(22740)

x+y =?
x-7 = 5
x*y = 24
:
it's easy to see from the 2nd equation that x = 12, then y = 2)
:
Assuming this is what you meant:
x+y = ?
x-y = 5
x*y = 24
:
Rearrange the 2nd equation for substitution
x = (y+5)
:
Substitute in the 3rd equation
y(y+5) = 24
y^2 + 5y - 24 = 0; a quadratic equation
Factor
(y + 8)(y - 3) = 0
Two solutions
y = 3, then x = 8
y = -8, then x = -3
:
Therefore x + y = ?
8 + 3 = 11
and
-3 + (-8) = -11